Operations Report – November 29th

rew 216 Operations Report 29-Nov-19
SOL: 5
Name of person filing report: Mike Lawson
Non-nominal systems: RAM rover power charging stations inoperable, electrical issue, ongoing.
Notes on non-nominal systems:
Generator: run
Hours run: 12.5
From what time last night: 1758
To what time this morning: 0630
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 71 %
Diesel Reading – 39
Station Propane Reading – 57
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1462858
Water (static tank) (gal): 343
Static to Loft Pump used – yes
Water in Green Hab (gal): 267
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: no
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging:
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: no
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100
Currently charging: yes
Notes on rovers: will have to juggle charging sequence.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments: Weather day, no EVA ativity. HAB maintenance, mission planning for future EVAs and personal project/report time.
Summary of internet: No serious issues.
Summary of suits and radios: No issues.
Summary of Hab operations: Routine maintenance and housekeeping.
Summary of GreenHab operations: Routine watering.
Summary of Science Dome operations: Gypsum processing continued.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved.
Summary of any observatory issues: None, weathered out.
Summary of health and safety issues: HSO broke sim to address medical issue at Base Sopport.
Questions, concerns and requests to Mission Support: Need to replace 1 bottle of cough suppresent/expectorant consumed.

Journalist Report – November 29th

Crew 216 Journalist Report 29-NOV-19
Sol 5
Author: Evgenia Alexandrova

Do you believe in miracles? Yes or no? If you can’t answer straight to that, choose between these two options: most likely yes or most likely no? What you answer actually describes how much you trust life. I think coming to MDRS was a choice of faith. In return today, we were gratified with a real space miracle: snow on Mars. It was bewitching and relaxing to watch the snowflakes smash the window in the science dome. Most of us find it cozy and enjoyable staying indoors and meditate while the weather gets angry outside. This is interesting, because almost each crewmember describes himself as an “outdoor person”. So how come a bunch of intelligent, communicative, OUTDOOR people decided free willingly to isolate themselves in a 75m2 Hab? The answer was found: Mars is definitely outdoor to Earth and to us it means adventure. This is the best incentive.

The freshest poem from an anonymous crew member:

As we wait and look outside

And watch the storm without

The mix of snow and rain abide

And dark clouds hang about

Research Report – November 29th

Crew 216 Research Report 29-Nov-2019
Sol 5
Submitted by: Crew Researcher Whittle

1) Psych Scale (TAMU IRB2019-1042D) – Completed second set of psychological surveys today.

2) Medical Casts – Continuation of Crew 215 research. Crushed 40% of gypsum, will continue tomorrow and begin purifying.

3) In-Situ Fluorescent Mineral Prospecting – Nothing further to report.

4) MDRS Mapping – Nothing further to report.

5) MDRS Communications – Completed GIS viewsheds of potential repeater sites to West and South of Hab.

Commander Report – November 29th

Crew 216 Commander Report 29-NOV-2019
Sol: 5
Summary Title: Quiet Day
Author’s name: Marc Levesque
Crew 216 spent the post-holiday hunkered down because of very wet and windy weather that included snow. Despite this, we were able to undertake a variety of activities related to household chores and projects, including Evgenia interviewing me for her film. I wasn’t sure how comfortable I would be, but I hardly notice the camera sitting in front of me, mostly because of her easygoing interviewing style. Some of her questions were quite thought provoking, forcing me to pause before answering, and I hope she was able to capture what she had hoped. The day also provided an opportunity for Rich to write some graduate school papers, while Michael spent time creating a spreadsheet for a GreenHab checklist. Andrew was busy in the Science Dome working on the gypsum samples, and Mike continues to think and work about so many things to help improve operations around the Hab. What a great crew. The day wound down with Mike and me working up some Martian chile with rice that the crew seemed to enjoy. We look forward to seeing some sun tomorrow.

GreenHab Report – November 29th

Crew 216 GreenHab Report 29-NOV-19
Crew GreenHab Officer: Michael Ho

Environmental control: Door opened for ventilation at 11.55 am, closed at 5.10 pm.
Average temperature: 14C
Hours of supplemental light: Automatic system
Daily water usage for crops: 12L
Daily water usage for research and/or other purposes: nil
Water in Blue Tank – enough
Time(s) of watering for crops: 09.40 am; 4.10 pm.

(Established plants introduced since 03/11/2019 till 22/11/2019 please refer to Crew 215 GreenHab Report of 22/11/2019.)

Constructed temporary shades for the 3 strawberry plants for shielding from heater vent. Will monitor over the coming days.

Rearranged on platform nearest to blue water container (Plot A):
a. the 9 pots of Tomato into 3×3 matrix,
b. the 4 pots of Sweet Chard into 2×2 matrix,
c. the 4 pots of Sweet Peppers into 2×2 matrix,
d. the 2 pots of Chile Pepper into 1×2 matrix,
e. the 2 pots of Avocado into 1×2 matrix.

Harvest: None
Support/supplies requested: None

Sol Summary – November 29th

Crew 216 Sol Summary Report 29-NOV-19
Sol: 5
Summary Title: Catch up day
Author’s name: Marc Levesque
Mission Status: Active
Sol Activity Summary: The crew took advantage of the windy and wet weather that included snow to catch up on indoor activities and more importantly, rest. Filming continued for Evgenia’s project, Andrew worked on the gypsum project, Michael spent time developing a GreenHab checklist, while others caught up on graduate research papers, completed household chores, or napped.
Look Ahead Plan: An EVA will be requested for tomorrow afternoon after the ground has a chance to dry out to check out the potential north repeater site that was not reached on Wednesday. We will also tend to some household chores in anticipation of the Roker film crew arriving on Sunday.
Anomalies in work: N/A.
Weather: Cold and wet overnight.
Crew Physical Status: Healthy
Reports to be filed: Sol Summary Report, Commander Report, Operations Report, GreenHab Report, Journalist Report, Research Report, and Photos.
Support Requested: None

Commander Report – November 27th

Crew 216 Commander Report 27-NOV-2019

Sol: 3
Summary Title: More EVAs
Author’s name: Marc Levesque
Crew 216 took advantage of the good weather to continue field work in support of its projects. Of particular note was searching for potential radio repeater sites in the hopes of improving communications between the Hab and EVA teams. Documentary and drone footage was also collected during these activities. The crew has fallen into a comfortable routine of life at MDRS, though inconsistent sleep seems to be elusive for many of the crew. The dryness of the Martian climate is a challenge for those who hail from more temperate and moist climes. Overall, however, the crew seems healthy with a couple of members on the backside of head colds picked up in their travels to the station. Note: Some crew members are unable to send their reports because they are unable to access their email server. As a result, the Crew Commander is sending those.

Astronomy Report – November 27th

Crew 216 Astronomy Report 27-NOV-19
Crew Astronomer: Michael Ho

Robotic Telescope Requested: none
Objects to be Imaged this Evening: none
Images submitted with this report: nil
Problems Encountered: nil

Solar Features Observed: No prominences observed.
Images submitted with this report: nil
Problems Encountered: There is a gap in the fence outside the observatory on the north side that I might have missed before. Is that normal?

EVA Report – November 27th

Crew 216 EVA Report 27-Nov-19
EVA # 05
Author: Levesque
Purpose of EVA: To explore suitability for possible radio repeater sites close to the Hab. Collect footage for documentary.
Start time for EVA: 0930
End time for EVA: 1130
Narrative: EVA #5 was our crew’s first non-rover EVA. We left the Airlock at 0930 with Levesque, Lawson, Alexandrova, and Kennedy. We traveled west up the bentonite hills adjacent to the Hab, stopping a few times for footage and oxygen absorption. At the summit of the ridge we were greeted by the majestic views in all directions, looking down onto our decorous Hab below. We then traveled south to the coordinate #1. We observed a surprising amount of fossils for Mars, but found the road in good condition and excellent for travel. From location #1 we stopped, did some filming, and returned north on the road we had been traveling. We arrived at the coordinate #2 shortly after. This is the site of the old repeater according to our documents. We evaluated the suitability of the location. It would do well for comms east west and south, but site #3 would most likely provide a better overall coverage, especially north. We did some more filming with only minor queues from our documentarian and continued north on the Hab Ridge road. Shortly after we came to the break in the ridge and the old access road. We followed this down until we were at the foot of the north ridge, the base of the hill of coordinate #3. We were concerned that with the time we had left that it would be cutting it close, so we elected to reevaluate the location at a later EVA and return to the Hab. We did some minor filming and savored how we are living our best lives. We did a quick external evaluation of the Hab and returned to the airlock approximately at 1130.
Destination: In the vicinity of Hab (within 1 km), 3 specific points listed below.
Coordinates (use UTM NAD27 CONUS):
1: 12 S 0517874 4250342.
2: 12 S 0517887 4250342.
3: 12 S 0518464 4251538.
EVA Participants: Levesque, Lawson, Alexandrova, Kennedy
Road(s) and routes per MDRS Map:
1) Hab to 12 S 0517874 4250342 on foot
2) 12 S 0517874 4250342 to 12 S 0517887 4250342 on foot
3) 12 S 0517887 4250342 to 12 S 0518464 4251538 on foot
4) 12 S 0518464 4251538 to Hab on foot
5) Total distance 3.5 km
Mode of travel: Walking with powerful self determination
Vehicles used: N/A

Copyright © The Mars Society. All rights reserved. | Main Site